OpenStudy (anonymous):

Find the derivative of the trigonometric function. y=2xsinx+x^2cosx

OpenStudy (diyadiya):

\[y'=\frac{d}{dx}(2xsinx)+ \frac{d}{dx}x^2cosx \]

OpenStudy (cwrw238):

use product rule

OpenStudy (anonymous):

product rule for addition?

OpenStudy (diyadiya):

use product rule for each term

OpenStudy (anonymous):

ah ok

OpenStudy (cwrw238):

no - apply product rule to each of the two terms - looks like diyad is doing that now

OpenStudy (diyadiya):

First do \( \frac{d}{dx}(2xsinx) \) & then \( \frac{d}{dx}(x^2cosx) \)

OpenStudy (diyadiya):

do you know how to do derivative of 2xsinx ?

OpenStudy (cwrw238):

d (uv)dx = u* dv/dx + v* du/dx where u and v are function of x

OpenStudy (anonymous):

so

OpenStudy (anonymous):

2xsix+x^2cosx=2sinx+3xcosx+2xcosx-x^2sinx

OpenStudy (anonymous):

2sinx oops

OpenStudy (diyadiya):

Nope

OpenStudy (anonymous):

2xsinx

OpenStudy (anonymous):

4xcosx+2sinx-x^2sinx

OpenStudy (anonymous):

4xcosx+sinx(2-x^2)

OpenStudy (anonymous):

i mistyped a bunch at first lol

OpenStudy (diyadiya):

\[\frac{d}{dx} (2xsinx)= 2 \frac{d}{dx}(xsinx) =2[x \frac{d}{dx}sinx+sinx \frac{d}{dx}x]\]

OpenStudy (anonymous):

my last answer is correct

OpenStudy (anonymous):

4xcosx+sinx(2-x^2)

OpenStudy (anonymous):

atleast thats what the books says in the back.

OpenStudy (diyadiya):

\[ \frac{d}{dx}(2xsinx)=2[xcosx+sinx]\]

OpenStudy (anonymous):

y=2xsinx+x^2cosx = 2sinx+2xcosx+2xcosx-x^2sinx = 4xcosx +2sinx - x^2sinx = 4xcosx + sinx(2-x^2)

OpenStudy (cwrw238):

thats correct ChrisV

OpenStudy (anonymous):

thanks guys

OpenStudy (diyadiya):

oh sorry i thought you were taling about the derivative of 2xsinx

OpenStudy (anonymous):

did not know i could use product rule on both sides

OpenStudy (anonymous):

better to learn than wonder :)

OpenStudy (cwrw238):

yep - whenever you have a product of two functions the rule applies